OpenStudy (sonic9678):

I'm not sure if I'm overthinking this question, so could someone please walk me through this? :): 'If the margin of error in an estimate for the mean weight of a shipment is + or -2 pounds at a confidence level of 95 percent, what will be the margin of error at a confidence level of 98 percent? Be sure to show how you arrived at your answer.' Thanks

OpenStudy (mww):

http://www.flashandmath.com/mathlets/statistics/meansci/ciformula2.jpg This formula will be applied for confidence interval range. (or use z* instead of t* if it is a population) So the margin of error is just z*s/sqrt(n)) Find z* (critical z value based on your tables), for 95% CI and then rearrange the equation to find s/sqrt(n) Then multiply this s/sqrt(n) by z* for 98% CI and you're done The point is the standard error, i.e. s/sqrt(n) does not change, but the z* given by confidence level does.

OpenStudy (michele_laino):

if the involved distribution is the gaussian distribution, then we can write: \[1.96 \cdot \sigma = 2\] wherein \(\sigma\) is the standard deviation please find \(\sigma\)

OpenStudy (sonic9678):

@mww ahh okay, that makes sense! Thank you so much :) also @Michele_Laino tysm

OpenStudy (michele_laino):

:) the requested uncerainty, is: \[2.33 \cdot \sigma = ...?\]

OpenStudy (agent0smith):

Margin of error is the z score times the standard error.\[ME = z* SE\]since you know the margin of error, and the z score for 95% confidence (1.96), you can find the Standard Error. Then just plug in the z score for 98% confidence (get it from a z table) and multiply by the same SE you just worked out.

OpenStudy (agent0smith):

Though it looks like you might want to use t* since i'm guessing it's a sample.

OpenStudy (sonic9678):

awesommme thank you all your explanations helped a lot !! @agent0smith we're using z :) ty!!